QUANTUM CHEMICAL CLUSTER STUDIES OF ICE-BOUND REACTIONS OF FORMALDEHYDE (H_{2}CO), ACETALDEHYDE (CH_{3}CHO), OR ACETONE (CH_{3}COCH_{3}) WITH AMMONIA (NH_{3})

\underline{DAVID E. WOON} and DOROTHY J. MILLER, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana IL 61801.

While many of the reactions that can occur in icy grain mantles in cold interstellar clouds are either barrierless recombinations of open-shell radicals or are driven by energetic processing (photolysis, radiolysis, or shocks), there is a group of unusual reactions involving stable, closed-shell species that can also be efficient at very low temperatures. Previous experimental work^{a} found that H_{2}CO and NH_{3} can evidentally react at temperatures under 70 K if they are embedded within an ice matrix, a conclusion that was subsequently confirmed theoretically^{b}. To assess the impact of including more water, cluster calculations were performed at the MP2/6-31+G** or B3LYP/6-31+G** level with up to 4H_{2}O and 12H_{2}O, respectively. Electrostatic contributions from bulk ice were modelled with either PCM and IPCM reaction field solvation. In addition to revisiting H_{2}CO + NH_{3}, we also characterized the reactions of ammonia with acetaldehyde and acetone. The new work confirms that the reaction barriers are very small, but it also indicates that water induces a substantial interaction between NH_{3} and the carbonyl species in the solvated pre-reaction complexes. In small clusters, the C-N distance decreases by about 1 Å, and partial charge-transfer complexes are formed. In large clusters, a proton transfers to yield cationic complexes. We will present an analysis of potentially observable vibrational features of these complexes.

aW. A. Schutte, L. J. Allamandola, and S. A. Sandford, Science 259, 1143 (1993); Icarus 104, 118 (1993).

bD. E. Woon, Icarus 142, 550 (1999).